OpenStudy (anonymous):

there are 3 caskets.1st one contains 2 gold coins, 2nd one contains 1 gold and 1 silver coins and 3rd one contains 2 silver coins.I choose one caskets at random and draw a coin.If the coin is gold,probability that the 2nd coin in the casket is gold can be written as a/b where a,b are co prime positive numbers.Then a+b=?

OpenStudy (shubhamsrg):

BRILLIANT.ORG !

OpenStudy (shubhamsrg):

:P

OpenStudy (shubhamsrg):

P(choosing 2nd casket) =1/3 P(gold coin from 2nd casket) = 1/2 P(gold from 1st) = 1 P(gold from 3rd) = 0 So required probab = (1/3)*(1/2) / [ (1/3)*1 + (1/3)*(1/2) + (1/3)*0 ] Just simplify this

OpenStudy (anonymous):

Can someone help me over modulo arithmatics